F. Determine the length of the parametric curve given by the following parametric equations. x=3sin(t) y=3cos(t) 0<t<2π


1
Expert's answer
2022-01-17T18:01:52-0500

Solution

For given parametric equations

"L=\int_{0}^{2\pi}{\sqrt{\left(\frac{dx}{dt}\right)^2+\left(\frac{dy}{dt}\right)^2}dt=}\int_{0}^{2\pi}{\sqrt{\left(3cos(t)\right)^2+\left(-3sin(t)\right)^2}dt=}"

"=3\int_{0}^{2\pi}{\sqrt{{cos(t)}^2+{sin(t)}^2}dt=}3\int_{0}^{2\pi}dt=6\pi"


Answer

L=6π